OpenStudy (anonymous):

45=7-5m (PLEASE SHOW WORK)

OpenStudy (anonymous):

45 = 7-5m 45 -7 = 7 -7 -5m 38 = -5m -5m/-5 = 38/-5 m = -38/5 or -7 1/5

OpenStudy (anonymous):

it says that the answer is supposed to be -7 3/5?